What are the numerical expression of...
5 decreased by the product of a number and 3?
The sum of -9 times a number and 6 is 72?
Thank-you for your consideration if you have time to answer my question.

Answer by Maths68(1474)

Let number = x
(1) 5 decreased by the product of a number and 3 = 3x-5

(2) The sum of -9 times a number and 6 is 72 = 6-9x=72